OpenStudy (anonymous):

Two numbers differ by 9. The sum of their squares is 653. Find the numbers.

OpenStudy (paxpolaris):

Let the two numbers be a and b \[a-b=9 \]\[a^2+b^2=653 \]

OpenStudy (anonymous):

And I should rearrange the a and the b to plug it into the second equation, right?

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

How do you factor 2b^2+188b-653?

OpenStudy (paxpolaris):

first we need to know factors of 653 ... can't be 3 or 11. .....try 7

OpenStudy (anonymous):

7 doesn't work

OpenStudy (paxpolaris):

13

OpenStudy (paxpolaris):

***how about 13 you could always use ... \[\Large x= {-b \pm \sqrt{b^2-4ac} \over 2a}\]

OpenStudy (anonymous):

Sorry, it's been about 2 years since I've used that formula. Can you just remind me how to do it? @PaxPolaris

OpenStudy (paxpolaris):

sorry.... didn't notice that the quadratic equation you have is actually Wrong. \[\large a^2+b^2=653\]\[\implies \large \left( b+9 \right)^2+b^2=653\ \dots\ \dots \left\{ \because a=b+9 \right\}\]\[\implies \Large \left( b^2+18b+81 \right)+b^2=653\]implies Large \left( b^2+18b+81 \right)+b^2=653

OpenStudy (paxpolaris):

\[\implies \Large 2b^2+18b+81 -653=0\]\[\implies \LARGE 2b^2+18b-572=0\] .... then finally you can divide everything by 2:\[\implies b^2+9b-286=0\]